Thanks for your help. Can anyone direct me on this one?
P(-3,-7); m= -2/3
Use slope-intercept form to write the equation of line passing through the given point and having the given slope. Express the answer in general form.
Thanks!

Answer by checkley71(8403)

Y=mX+b Y=-7,X=-3 m=-2/3 THUS
-7=-2/3*-3+b
-7=6/3+b
-7=2+b
b=-7-2
b=-9
THUS THE EQUATION IN STANDARD FORM IS
Y=-2/3X-9